# The Endless Tuning An Artificial Intelligence Design To Avoid Human Replacement and Trace Back Responsibilities

### v1

@@ 4.2 Insights from user experience
With regard to the evaluation of the interface, we observed some intriguing discrepancy between the interviewee’s responses and his actual behavior. When asked about the usefulness of the rule-based suggestions (see Figure 3(a)), the interviewee acknowledged their relevance, particularly in relation to a comparison between the “loan percent income” threshold derived from the decision tree and that applied by the bank. However, it is worth noting that this aspect became a focal point of the discussion because, while reading the suggestions, the interviewee misread a value and was corrected by the interviewer («Oh right — our bank’s threshold is 33–35%, that’s why I got confused»). On the one hand, this may suggest the presence of an anchoring bias (see again [4]), which could indeed be mitigated; on the other hand, it reveals the need for a more distinct and prominent interface than the brightly colored text currently employed. Concerning the similarity, the interviewee answered skeptically («because I evaluate each case individually»). While it must be noted that — both in this and in the other two cases — the chart proved to be of little use (perhaps, we may hypothesize, because it displayed relative distances without providing a fixed point of reference), it is nonetheless noteworthy that, despite his own statement, the director spent a considerable amount of time carefully examining the similar cases (see Figure 3(b)). Without suggesting that the interviewee was in any way unaware, this observation seems to warrant further investigation aimed at catching user attention.
Figure 4: (a) The saliency map produced with the Grad-CAM [89, 76] explainer, regarding the picture chosen in the art style recognition test. (b) The corresponding RISE [75] explanation (both maps presented before knowing the outcome, exploring a hermeneutic strength). (c) The picture chosen by the professor of aesthetics. Notably, it is a painting of 1659 by Francisco de Zurbarán representing Saint Francis.
@@
Figure 6: (a) The saliency map produced with the Grad-CAM [89, 76] explainer, regarding the picture chosen in the pneumonia diagnosis test. (b) The corresponding RISE [75] explanation (both maps presented before knowing the outcome, exploring a hermeneutic strength). (c) The chest x-ray picture chosen by the radiologist.
@@ 5 Filling gaps
Such a design possesses a somewhat vertical character and appears well suited to medium-to-high-stake situations. Our aim is to place emphasis on the trust granted to the user for what they are, through affordances too, continuously re-building trust itself or perhaps, using a more concrete and engaging term, enhancing reliability. A valid objection could be that such a relationship must be oriented toward improvement, avoiding the decay of information. Now, if an artificial intelligence system manages to provide useful clues that are meaningful in their own way, such information cannot be elevated to the status of knowledge. Strictly speaking, this, to us, is the ethics of artificial intelligence. The genitive is objective, not subjective. One should contribute something of their own: a computer program alone cannot solve all our problems and yield honest and serious and competent people out of us. As previously mentioned, a slightly larger margin of error might be tolerated — whereas, statistically speaking, it would otherwise always be deemed as a fortuitous case — if it would be governable: the interviewees reported full control. Broadly speaking, an approach oriented to ethics of care seems to emerge, where principles certainly matter but, without underestimating the centrality of fundamental rights, it is acknowledged that «the morality of rights differs from the morality of responsibility in its emphasis on separation rather than connection, in its consideration of the individual rather than the relationship as primary»[3, p. 19]. In short, ethics is conceived as a field with a variable center of gravity. In the beginning is the relation [16]. It is primarily the operator and, more or less directly, the dataset annotators, developers, interface designers, etc. who relate to a potentially harmed party, except for the reversal of the burden of proof by the entire supply chain. As a consequence, time plays a role: responsibility is created “live”, in the very acting, and it is a responsibility of using the device but also for the people (note, in passing, that the product’s flaw has itself a relational nature [18, pp. 240 ff.]). On the other hand, it does not simply hold that «the journey prevails over the object» [90, p. 153] [^5] Translation is mine.. Such temporality might be perhaps described as a “delayed virtuality”.
@@
Such a project seems to align with [91] when, concerning a computational accountability to trigger a «multilayer accountability system»(also referring to a definition of accountability from a computational perspective by the IEEE [92, p. 3]), it contends to be possible «for computing itself to enable the almost automatic verification of the causal reasons behind AI actions»[p. 1011] [^6] Translation is mine., so that the “bill” could be effectively distributed. Surely, the AI Act fits this approach, providing precise obligations towards safety [18, p. 360], but a link exists too to the other face of responsibility, i.e. liability. As an example, the core of the AI Liability Directive proposal, actually retired by the EU Commission in 2025777See https://www.ansa.it/osservatorio_intelligenza_artificiale/notizie/progetto_solaris/2025/02/12/la-commissione-ue-ritira-la-direttiva-sulla-responsabilita-da-ia_00ae0807-d41d-4ed5-a225-1d938d27af4b.html (accessed 06/06/2025)., nonetheless consisted of providing «an economic incentive to comply with safety rules», contributing «to the enforcement of the requirements for high-risk AI systems imposed by the AI Act, because the failure to comply with those requirements constitutes an important element triggering the alleviations of the burden of proof»[93]. Nevertheless, it could perhaps be asked how straight such a link is and, again, if the structure itself of the software could not help bridging these two banks, up to touching (at the extreme side of a spectrum starting from guilt or alleged guilt) objective liability, maybe facilitating the application of the case in point of damages provoked by goods held in trust (which, at least w.r.t. the Italian Civil Code (art. 2051), seems the most appropriate [18, p. 148]).
@@
A positive note might instead regard the recital 53 of the AI Act [19, pp. 14-15], concerning «specific cases in which AI systems referred to in pre-defined areas specified in this Regulation do not lead to a significant risk of harm to the legal interests protected under those areas». Among the conditions that should be fulfilled in order for a system to take a step down in the pyramid of risk, for example, the second one states that «the task performed by the AI system is intended to improve the result of a previously completed human activity that may be relevant for the purposes of the high-risk uses […]. Considering those characteristics, the AI system provides only an additional layer to a human activity with consequently lowered risk». We believe that the Endless Tuning could align at least with such requirement, perhaps entailing not insignificant advantages also from an economic viewpoint.

### v3

@@ 4.2 Insights from user experience
With regard to the evaluation of the interface, we observed some intriguing discrepancy between the interviewee’s responses and his actual behavior. When asked about the usefulness of the rule-based suggestions (see Figure 3(a)), the interviewee acknowledged their relevance, particularly in relation to a comparison between the “loan percent income” threshold derived from the decision tree and that applied by the bank. However, it is worth noting that this aspect became a focal point of the discussion because, while reading the suggestions, the interviewee misread a value and was corrected by the interviewer («Oh right — our bank’s threshold is 33–35%, that’s why I got confused»). On the one hand, this may suggest the presence of an anchoring bias (see again [kahneman2011thinking]), which could indeed be mitigated; on the other hand, it reveals the need for a more distinct and prominent interface than the brightly colored text currently employed. Concerning the similarity, the interviewee answered skeptically («because I evaluate each case individually»). While it must be noted that — both in this and in the other two cases — the chart proved to be of little use (perhaps, we may hypothesize, because it displayed relative distances without providing a fixed point of reference), it is nonetheless noteworthy that, despite his own statement, the director spent a considerable amount of time carefully examining the similar cases (see Figure 3(b)). Without suggesting that the interviewee was in any way unaware, this observation seems to warrant further investigation aimed at catching user attention.
Figure 4: (a) The saliency map produced with the Grad-CAM [jacobgilpytorchcam, selvaraju2017grad] explainer, regarding the picture chosen in the art style recognition test. (b) The corresponding RISE [petsiuk2018rise] explanation (both maps presented before knowing the outcome, exploring a hermeneutic strength). (c) The picture chosen by the professor of aesthetics. Notably, it is a painting of 1659 by Francisco de Zurbarán representing Saint Francis.
@@
Figure 6: (a) The saliency map produced with the Grad-CAM [jacobgilpytorchcam, selvaraju2017grad] explainer, regarding the picture chosen in the pneumonia diagnosis test. (b) The corresponding RISE [petsiuk2018rise] explanation (both maps presented before knowing the outcome, exploring a hermeneutic strength). (c) The chest x-ray picture chosen by the radiologist.
@@ 5 Filling gaps
Such a design possesses a somewhat vertical character and appears well suited to medium-to-high-stake situations. Our aim is to place emphasis on the trust granted to the user for what they are, through affordances too, continuously re-building trust itself or perhaps, using a more concrete and engaging term, enhancing reliability. A valid objection could be that such a relationship must be oriented toward improvement, avoiding the decay of information. Now, if an artificial intelligence system manages to provide useful clues that are meaningful in their own way, such information cannot be elevated to the status of knowledge. Strictly speaking, this, to us, is the ethics of artificial intelligence. The genitive is objective, not subjective. One should contribute something of their own: a computer program alone cannot solve all our problems and yield honest and serious and competent people out of us. As previously mentioned, a slightly larger margin of error might be tolerated — whereas, statistically speaking, it would otherwise always be deemed as a fortuitous case — if it would be governable: the interviewees reported full control. Broadly speaking, an approach oriented to ethics of care seems to emerge, where principles certainly matter but, without underestimating the centrality of fundamental rights, it is acknowledged that «the morality of rights differs from the morality of responsibility in its emphasis on separation rather than connection, in its consideration of the individual rather than the relationship as primary»[gilligan1993different, p. 19]. In short, ethics is conceived as a field with a variable center of gravity. In the beginning is the relation [Buber1970-BUBIAT]. It is primarily the operator and, more or less directly, the dataset annotators, developers, interface designers, etc. who relate to a potentially harmed party, except for the reversal of the burden of proof by the entire supply chain. As a consequence, time plays a role: responsibility is created “live”, in the very acting, and it is a responsibility of using the device but also for the people (note, in passing, that the product’s flaw has itself a relational nature [dal2024intelligenza, pp. 240 ff.]). On the other hand, it does not simply hold that «the journey prevails over the object» [virilio1989macchina, p. 153] [^5] Translation is mine.. Such temporality might be perhaps described as a “delayed virtuality”.
@@
Such a project seems to align with [comande2018responsabilita] when, concerning a computational accountability to trigger a «multilayer accountability system»(also referring to a definition of accountability from a computational perspective by the IEEE [ieee_glossary2017, p. 3]), it contends to be possible «for computing itself to enable the almost automatic verification of the causal reasons behind AI actions»[p. 1011] [^6] Translation is mine., so that the “bill” could be effectively distributed. Surely, the AI Act fits this approach, providing precise obligations towards safety [dal2024intelligenza, p. 360], but a link exists too to the other face of responsibility, i.e. liability. As an example, the core of the AI Liability Directive proposal, actually retired by the EU Commission in 2025, nonetheless consisted of providing «an economic incentive to comply with safety rules», contributing «to the enforcement of the requirements for high-risk AI systems imposed by the AI Act, because the failure to comply with those requirements constitutes an important element triggering the alleviations of the burden of proof»[eu_com2022_ai_liability]. Nevertheless, it could perhaps be asked how straight such a link is and, again, if the structure itself of the software could not help bridging these two banks, up to touching (at the extreme side of a spectrum starting from guilt or alleged guilt) objective liability, maybe facilitating the application of the case in point of damages provoked by goods held in trust (which, at least w.r.t. the Italian Civil Code (art. 2051), seems the most appropriate [dal2024intelligenza, p. 148]).
@@
A positive note might instead regard the recital 53 of the AI Act [EU2024AIAct, pp. 14-15], concerning «specific cases in which AI systems referred to in pre-defined areas specified in this Regulation do not lead to a significant risk of harm to the legal interests protected under those areas». Among the conditions that should be fulfilled in order for a system to take a step down in the pyramid of risk, for example, the second one states that «the task performed by the AI system is intended to improve the result of a previously completed human activity that may be relevant for the purposes of the high-risk uses […]. Considering those characteristics, the AI system provides only an additional layer to a human activity with consequently lowered risk». We believe that the Endless Tuning could align at least with such requirement, perhaps entailing not insignificant advantages also from an economic viewpoint.